OpenStudy (anonymous):

BOOYAHKAH BOOOOYAHKAH I NEED HELP PLEASE! If 1/x + 1/3 + 1/4 = 1, then x =

hero (hero):

First add 1/3 + 1/4

OpenStudy (anonymous):

4/12 3/12

OpenStudy (anonymous):

now what

hero (hero):

Add 4/12 + 3/12

OpenStudy (anonymous):

7/12

OpenStudy (kirbykirby):

\[\frac{1}{x}+\frac{7}{12}=1\\ \\ \, \\\frac{1}{x}=1-\frac{7}{12} \\ \, \\ \frac{1}{x}=\frac{5}{12} \text{ since }1=\frac{12}{12} \] Cross multiply: \[5x = 12\\ x = \frac{12}{5}=2.4\]

hero (hero):

I guess @kirbykirby finished it. @yomamabf do you understand it or have any questions?